# Variational Search Distributions

### v2

@@ 4.2 Fitness Landscapes
In this setting we wish to find $\mathbf{x}\in\mathcal{S}$, so we fix $\tau$ over all rounds for all competing methods. We only consider the combinatorially (near) complete datasets to avoid any pathological behavior from relying on machine learning oracles (Surana et al., 2024). Results are presented in Figure 3. The primary measures by which we compare methods are precision, recall and performance (the last adapted from Jain et al. (2022)),
$\textrm{Precision}_{t}$ $=\frac{1}{\min\{tB,|\mathcal{S}|\}}\sum^{t}_{r=1}\sum^{B}_{b=1}\mathbbm{1}[{y_{br}>\tau}]\cdot\mathbbm{1}[{\mathbf{x}_{br}\notin\mathcal{X}^{q}_{b-1,r}}],$ (15)
$\textrm{Recall}_{t}$ $=\frac{1}{\min\{TB,|\mathcal{S}|\}}\sum^{t}_{r=1}\sum^{B}_{b=1}\mathbbm{1}[{y_{br}>\tau}]\cdot\mathbbm{1}[{\mathbf{x}_{br}\notin\mathcal{X}^{q}_{b-1,r}}],$ (16)
$\textrm{Performance}_{t}$ $=\sum^{t}_{r=1}\sum^{B}_{b=1}y_{br}\cdot\mathbbm{1}[{\mathbf{x}_{br}\notin\mathcal{X}^{q}_{b-1,r}}].$ (17)
Here $\mathcal{X}^{q}_{br}\subset\mathcal{X}$ is the set of experimentally queried sequences by the $b$th batch member of the $r$th round, including the initial training set. These measures are comparable among probabilistic and non probabilistic methods. Precision and recall measure the ability of a method to efficiently explore $\mathcal{S}$, where $\min\{tB,|\mathcal{S}|\}$ is the size of the selected set at round $t$ (bounded by the number of good solutions), and $\min\{TB,|\mathcal{S}|\}$ is the number of positive elements possible in the experimental budget. Strictly, recall should be normalized by $|\mathcal{S}|$, but we use $TB$ here since it may not be realistic to have the experimental budget to fully explore $\mathcal{S}$. Performance measures the cumulative fitness of the unique batch members, and unlike Jain et al. (2022) we do not normalize this measure.
For the DHFR and TrpB experiments we set maximum fitness in the training dataset to be that of the wild type, and $\tau$ to be slightly below the wild type fitness value (so we have $\sim\!10$ positive examples to train the CPE with). We use a randomly selected $N_{\textrm{train}}=2000$ below the wild-type fitness to initially train the CPE, we also explicitly include the wild-type. The thresholds and wild-type fitness values are; DHRF: $\tau=-0.1$, $y_{\textrm{wt}}=0$, TrpB: $\tau=0.35$, $y_{\textrm{wt}}=0.409$. We follow the same procedure for the TFBIND8 experiment, however, there is no notion of a wild-type sequence in this data, and so we set $\tau=0.75$, and $y_{\textrm{train max}}=0.85$. We use a uniform prior over sequences, $p({\mathbf{x}})=\prod^{M}_{m=1}\mathrm{Categ}({x_{m}}|{\mathbf{1}\cdot|\mathcal{V}|^{-1}})$, since these are relatively small search spaces, and the sub-sequences of nucleic/amino acids have been specifically selected for their task. Similarly, we find that relatively simple independent (mean-field) variational distributions of the form in Equation 20 and MLP based CPE s work best for these experiments (details in Sec. B.2).
@@
Figure 3: Fitness landscape results. Precision (Equation 15), recall (Equation 16) and performance (Equation 17) – higher is better – for the combinatorially (near) complete datasets, DHFR and TrpB and TFBIND8. The random method is implemented by drawing $B$ samples uniformly.
Results are presented in Figure 3. VSD is the best performing method by most of the measures, with the related method CbAS also often performing well. We have found the AdaLead and PEX evolutionary-search based methods to be effective on lower-dimensional problems (TFBIND8 being the lowest here), however we consistently observe their performance degrading as the dimension of the problem increases. We suspect this is a direct consequence of their random mutation strategies being suited to exploration in low dimensions, but less efficient in higher dimensions compared to the learned generative models employed by VSD, CbAS, and DbAS. Our modified version of BORE (which is just the expected log-likelihood component of Equation 11) performs badly in all cases, and this is a direct consequence of its variational distribution collapsing to a point mass. In a non-batch setting this behavior is not problematic, but shows the importance of the KL divergence of VSD in this batch setting. We replicate these experiments in Sec. C.1 using GP-PI, also backed by our guarantees. In all cases VSD’s results remain similar or improve slightly, whereas the other methods results remain similar or degrade. We also report on batch diversity scores in Sec. C.2.
@@ 4.3 Black-Box Optimization
In this experiment we aim to find the global maximizers, $(y^{*}\!,\mathbf{x}^{*})$, of the black-box function, ${f\!\centerdot}$. For this, we set $\tau_{t}$ adaptively by specifying it as an empirical quantile of the observed target values,
@@
where $\tilde{Q}^{t}_{y}$ is the empirical quantile function of targets at round $t$, $p_{t-1}$ is a percentile from the previous round, and $\eta\in[0,1]$ is a parameter that controls an annealing-like schedule for $\tau_{t}$ that prioritizes exploration of the fitness landscape in earlier rounds and exploitation of known fit regions in later rounds. This is a strategy loosely-similar to Srinivas et al. (2010). The measure of interest for these experiments is simple/instantaneous regret $\mathrm{r}_{t}$ which quantifies how close the methods get to finding the globally fittest sequence,
$\mathrm{r}_{t}=y^{*}-\max_{y}\{y_{bi}\}^{B,t}_{b=1,i=1},$ (19)
where $y^{*}$ is the fitness value of the fittest sequence $\mathbf{x}^{*}$. We use the higher dimensional AAV ($y^{*}{=}19.54$) and GFP ($y^{*}{=}4.12$) datasets to show that VSD can scale to higher dimensional problems. However, the $\mathcal{X}$ of these experiments is completely intractable to fully explore experimentally, and so we use a predictive oracle trained on all of the original experimental data as the ground-truth black-box function. This is the same strategy used in Brookes et al. (2019); Jain et al. (2022); Trabucco et al. (2021); Kirjner et al. (2024) among others, and we use the exact CNN-based oracles from Kirjner et al. (2024) for these experiments. However, we note here that some of the oracles used in these experiments do not predict well out-of-distribution (Surana et al., 2024), which limits their real-world applicability.
We follow Kirjner et al. (2024) in the experimental settings for the AAV and GFP datasets, but we modify the maximum fitness training point and training dataset sizes to make them more amenable to a sequential optimization setting. The initial percentiles, schedule, and max training fitness values are; AAV: $p_{0}=0.8$, $\eta=0.7$, $y_{\textrm{max}}=5$, GFP: $p_{0}=0.8$, $\eta=0.7$ $y_{\textrm{max}}=1.9$. The edit distance between $\mathbf{x}^{*}$ and the fittest sequence in the CPE training data is 8 for GFP, and 13 for AAV. We again use a random $N_{\textrm{train}}=2000$ for training the CPE s, which in this case are CNNs – architecture specifics are in Sec. B.2.
@@
Figure 4: BBO results. Simple regret (Equation 19) – lower is better – on GFP and AAV with independent and auto-regressive variational distributions. The PEX and AdaLead results are replicated between the plots, since they are unaffected by choice of variational distribution. The random method is implemented by drawing $B$ samples from the prior models.
In these higher dimensional settings, we find that performance of the methods heavily relies on using an informed prior (in the case of VSD and CbAS), or initial variational distribution (in the case of DbAS and BORE). To this end, we follow Brookes et al. (2019) and fit the initial variational distribution to the CPE training sequences (regardless of fitness), but we use maximum likelihood. Then for VSD and CbAS we copy this distribution and fix its parameters for the remainder of the experiment for use as a prior. We also use this prior for the Random method, and AdaLead and PEX use alternative generative heuristics. For these experiments we use the simple independent variational distribution from before, and also the same LSTM and causal decoder-only transformer models from Sec. 4.1. We present additional results in Sec. C.3 with uninformed priors and transition variational distributions of the form $q({\mathbf{x}_{t}}|{\mathbf{x}_{t-1},\phi})$.
The results are summarized in Figure 4. VSD is among the leading methods for all experiments. VSD appears better take advantage of the more complex variational distributions than the competing methods CbAS and DbAS since it can sample from the adapted variational distribution while learning it. We can see that AdaLead, PEX and often BORE all perform worse than random for reasons previously mentioned. Simple regret can drop below zero for these experiments since an oracle is used as the black box function, but the global maximizer is taken from the experimental data. This potentially highlights some of the overconfidence issues when using oracles as outlined in Surana et al. (2024). Batch diversity scores for these experiments are presented in Sec. C.2.
@@ 5 Conclusion
We have presented variational search distributions (VSD), a method for efficiently finding designs of a rare class sequentially under some experimental constraints. VSD is underpinned by variational inference, which allows it to satisfy several critical requirements and important desiderata specific to this problem, including learning generative models for fit sequences and batch candidate generation. We show that VSD converges asymptotically to the true level-set distribution at the same rate as a Monte-Carlo estimator with full knowledge of the true distribution. We have also showcased the benefits of our method empirically on a set of combinatorially complete and high dimensional sequential-design biological problems and show that it can effectively learn powerful generative models of fit designs. Finally, our framework can be generalized to more complex application scenarios, potentially involving other challenging combinatorial optimization problems (Bengio et al., 2021), such as graph structures (Annadani et al., 2023), and mixed discrete-continuous variables, which are worth investigating as future work directions.
@@ Appendix A Acronyms
Bayesian optimization with probabilistic reparameterisation
@@
Appendix B Experimental Details
@@
We use three well established datasets; a green fluorescent protein (GFP) from Aequorea Victoria (Sarkisyan et al., 2016), an adeno-associated virus (AAV) Bryant et al. (2021); and DNA binding activity to a human transcription factor (TFBIND8) (Trabucco et al., 2022; Barrera et al., 2016). These datasets have been used variously by Brookes & Listgarten (2018); Brookes et al. (2019); Angermueller et al. (2019); Kirjner et al. (2024); Jain et al. (2022) among others. The GFP task is to maximize fluorescence, this protein consists of 238 amino acids, of which 237 can mutate. The AAV task us to maximize the genetic payload that can be delivered, and the associated protein has 28 amino acids, all of which can mutate. A complete combinatorial assessment is infeasible for these tasks, and so we use the convolution neural network oracle presented in Kirjner et al. (2024) as in-silico ground truth. TFBIND8 contains a complete combinatorial assessment of the effect of changing 8 nucleotides on binding to human transcription factor SIX6 REF R1 (Barrera et al., 2016). The dataset we use contains all 65536 sequences prepared by Trabucco et al. (2022). We also use two novel datasets from recent works that experimentally assess the (near) complete combinatorial space of short sequences. The first dataset measures the antibiotic resistance of Escherichia coli metabolic gene folA, which encodes dihydrofolate reductase (DHFR) (Papkou et al., 2023). Only a sub-sequence of this gene is varied (9 nucleic acids which encode 3 amino acids), and so a near-complete (99.7%) combinatorial scan is available. For variants that have no fitness (resistance) data available, we give a score of $-1$. The next dataset is near-complete combinatorial scan of four interacting amino acid residues near the active site of the enzyme tryptophan synthase (TrpB) (Johnston et al., 2024), with 159,129 unique sequences and fitness values, we use $-0.2$ for the missing fitness values (we do not use the authors’ imputed values). These residues are explicitly shown to exhibit epistasis – or non-additive effects on catalytic function – which makes navigating this landscape a more interesting challenge from an optimization perspective. The properties of these datasets are presented in Table 3.
@@
Table 3: Alphabet size, sequence length, and number of availbale sequences for each of the datasets we use in this work.
B.1 Variational distributions
In this section we summarize the main variational distribution architectures considered for VSD, BORE, CbAS and DbAS, and the sampling distributions for the Random baseline method. Somewhat surprisingly, we find that we obtain consistently good results for the biological sequence experiments using a simple independent (or mean-field) variational distribution,

### v4

@@ 4.2 Fitness Landscapes
In this setting we wish to find fit sequences $\mathbf{x}\in\mathcal{S}_{\textrm{SLS}}$, so we fix $\tau$ over all rounds. We only consider the combinatorially (near) complete datasets to avoid any pathological behavior from relying on machine learning oracles (Surana et al., 2024). Results are presented in Figure 3. The primary measures by which we compare methods are precision, recall and performance,
$\textrm{Precision}_{t}$ $=\frac{1}{\min\{tB,|\mathcal{S}|\}}\sum^{t}_{r=1}\sum^{B}_{b=1}\mathds{1}[{y_{br}>\tau}]\cdot\mathds{1}[{\mathbf{x}_{br}\notin\mathcal{X}^{q}_{b-1,r}}],$ (15)
$\textrm{Recall}_{t}$ $=\frac{1}{\min\{TB,|\mathcal{S}|\}}\sum^{t}_{r=1}\sum^{B}_{b=1}\mathds{1}[{y_{br}>\tau}]\cdot\mathds{1}[{\mathbf{x}_{br}\notin\mathcal{X}^{q}_{b-1,r}}],$ (16)
$\textrm{Performance}_{t}$ $=\sum^{t}_{r=1}\sum^{B}_{b=1}y_{br}\cdot\mathds{1}[{\mathbf{x}_{br}\notin\mathcal{X}^{q}_{b-1,r}}].$ (17)
Here $\mathcal{X}^{q}_{br}\subset\mathcal{X}$ is the set of experimentally queried sequences by the $b$th batch member of the $r$th round, including the initial training set. These measures are comparable among probabilistic and non probabilistic methods. Precision and recall measure the ability of a method to efficiently explore $\mathcal{S}$, where $\min\{tB,|\mathcal{S}|\}$ is the size of the selected set at round $t$ (bounded by the number of good solutions), and $\min\{TB,|\mathcal{S}|\}$ is the number of positive elements possible in the experimental budget. Performance measures the cumulative fitness of the unique batch members, but unlike Jain et al. (2022) we do not normalize this measure.
@@
Figure 3: Fitness landscape results. Precision (Equation 15), recall (Equation 16) and performance (Equation 17) – higher is better – for the combinatorially (near) complete datasets, DHFR, TrpB and TFBIND8. The random method is implemented by drawing $B$ samples uniformly.
For exact experimental settings we refer the reader to Appendix C.1. We set $\tau$ to be that of the wild-type sequences in the DHFR and TrpB datasets, and use $\tau=0.75$ for TFBIND8. We find that a uniform prior over sequences, and a mean field variational distribution (Equation 20) are adequate for these experiments, as is a simple MLP for the CPE. Results are presented in Figure 3. VSD is the best performing method by most of the measures. We have found the AdaLead and PEX evolutionary-search based methods to be effective on lower-dimensional problems (TFBIND8 being the lowest here), however we consistently observe their performance degrading as the dimension of the problem increases. We suspect this is a direct consequence of their random mutation strategies being suited to exploration in low dimensions, but less efficient in higher dimensions compared to the learned generative models employed by VSD, CbAS, and DbAS. Our modified version of BORE (which is just the expected log-likelihood component of Equation 11) performs badly in all cases, and this is a direct consequence of its variational distribution collapsing to a point mass. In a non-batch setting this behavior is not problematic, but shows the importance of the KL divergence of VSD in this batch setting. We replicate these experiments in Appendix D.1 using GP-PI, also backed by our guarantees. In all cases VSD’s results remain similar or improve slightly, whereas the other methods results remain similar or degrade. We report on batch diversity scores in Appendix D.3.
@@ 4.3 Black-Box Optimization
In this experiment we use VSD on the related task of BBO for finding $\mathcal{S}_{\textrm{BBO}}$. We set $\tau_{t}$ adaptively by specifying it as an empirical quantile $\tilde{Q}^{t}_{y}$ of the observed target values at round $t$,
@@
where $p_{t-1}$ is a percentile from the previous round, and $\eta\in[0,1]$ is an annealing parameter for $\tau_{t}$ that trades off exploration and exploitation. Performance is measured by simple regret $\mathrm{r}_{t}$, which quantifies the fitness gap between the globally optimal design and the best design found,
$\mathrm{r}_{t}=y^{*}-\max_{y}\{y_{bi}\}^{B,t}_{b=1,i=1}.$ (19)
Here $y^{*}$ is the fitness value of the globally optimal sequence $\mathbf{x}^{*}$. We use the higher dimensional AAV ($y^{*}{=}19.54$), GFP ($y^{*}{=}4.12$) and Ehrlich functions ($y^{*}{=}1$) datasets/benchmarks to show that VSD can scale to higher dimensional problems. $\mathcal{X}$ of AAV and GFP is completely intractable to fully explore experimentally, and so we use a predictive oracle trained on all the original experimental data as the ground-truth black-box function. We use the CNN-based oracles from Kirjner et al. (2024) for these experiments. However, we note here that some oracles used in these experiments do not predict well out-of-distribution (Surana et al., 2024), which limits their real-world applicability. The Ehrlich functions (Stanton et al., 2024) are challenging biologically inspired closed-form simulations that cover all $\mathcal{X}$. We compare against a genetic algorithm (GA), CbAS and LaMBO-2 (Gruver et al., 2023) for sequences of length $M=\{15,32,64\}$ using the poli and poli-baselines benchmarks and baselines software (González-Duque et al., 2024). For these experiments we use CNNs for the CPE s – all experimental settings are in Appendix C.2.
@@
Figure 4: AAV & GFP BBO results. Simple regret (Equation 19) – lower is better – on GFP and AAV with independent and auto-regressive variational distributions. The PEX and AdaLead results are replicated between the plots, since they are unaffected by choice of variational distribution.
(a) $M=15$
(b) $M=32$
(c) $M=64$
Figure 5: Ehrlich function (poli implementation) BBO results. VSD and CbAS with different variational distributions; mean field (MF), LSTM and transformer (TFM), compared against genetic algorithm (GA) and LaMBO-2 baselines.
The results are summarized in Figure 4 and 5. Batch diversity scores for these experiments are presented in Appendix D.3, and for holo Ehrlich function implementation results see Appendix D.2. VSD is among the leading methods for all experiments. VSD takes better advantage of the more complex variational distributions than CbAS and DbAS since it can sample from the adapted variational distribution while learning it. We can see that AdaLead, PEX and often BORE all perform worse than random for reasons previously mentioned. Simple regret can drop below zero for AAV & GFP since an oracle is used as the black box function, but the global maximizer is taken from the experimental data. VSD outperforms CbAS on the Ehrlich function benchmarks, and is competitive with LaMBO-2. We also present an ablation study in Appendix D.4.
@@ 5 Conclusion
We have presented the problem of active generation — sequentially learning a generative model for designs of a rare class by efficiently evaluating a black-box function — and a method for efficiently generating samples which we call variational search distributions (VSD). Underpinned by variational inference, VSD satisfies critical requirements and important desiderata, and we show that VSD converges asymptotically to the true level-set distribution at the same rate as a Monte-Carlo estimator with full knowledge of the true distribution. We showcased the benefits of our method empirically on a set of combinatorially complete and high dimensional sequential-design biological problems and show that it can effectively learn powerful generative models of fit designs. There is a close connection between active generation and black-box optimization, and with the advent of powerful generative models we hope that our explicit framing of generation of fit sequences will lead to further study of this connection. Finally, our framework can be generalized to more complex application scenarios, involving learning generative models over Pareto sets, $\mathcal{S}_{\textrm{Pareto}}$, in a multi-objective setting, or other challenging combinatorial optimization problems (Bengio et al., 2021), such as graph structures (Annadani et al., 2023), and mixed discrete-continuous variables. All of which are worth investigating as future work directions. For the code implementing the models and experiments in this paper, please see https://github.com/csiro-funml/variationalsearch.
Acknowledgments
This work is funded by the CSIRO Science Digital and Advanced Engineering Biology Future Science Platforms, and was supported by resources and expertise provided by CSIRO IMT Scientific Computing. We would like to thank the anonymous reviewers, and especially reviewer xamp for the constructive feedback and advice, which greatly increased the relevance and quality of this work.
@@ Appendix A Acronyms
Bayesian optimization with probabilistic reparametrization
@@
Appendix B Depiction of Active Generation
See Figure B.1 for graphical depictions of active generation as implemented by VSD, compared to batch BO. Active generation using VSD follows Algorithm 1 to sequentially approximate $p({\mathbf{x}}|{y>\tau})$. It uses samples from the current learned approximation of this distribution, $q({\mathbf{x}}|{\phi^{*}_{t}})$, for proposing candidates to evaluate each round. Unmodified batch BO in the discrete setting without any specialization requires a list of candidates, from which a batch of candidates are selected per round using a surrogate model with a batch acquisition function, e.g. see Wilson et al. (2017). The surrogate model’s hyper-parameters, $\theta$, are estimated by minimizing negative log marginal likelihood, $\mathcal{L}_{\textrm{NLML}}(\theta,\mathcal{D}_{N})$. Mechanistically, active generation learns a generative model of valuable candidates to circumvent the requirement of having to select candidates from a list. This is especially important for searching over the space of sequences, $\mathcal{X}$, where enumerating feasible candidates is often intractable. Furthermore, active generation naturally lends itself to large batch sizes, while explicit batch optimization is often computationally intensive and is limited to ${\mathcal{{O}}}(10)$ candidates per batch Wilson et al. (2017). As mentioned in section 3, alternatives to active generation for specializing batch BO to the discrete domain also include LSO (Tripp et al., 2020) and amortized BO (Swersky et al., 2020) among others.
(a) VSD Active Generation
(b) Batch BO
Figure B.1: Depictions of (1(a)) active generation as implemented by VSD, and (1(b)) batch Bayesian optimization as applied to discrete sequences. Please see the text for a discussion of the differences between these approaches.
Appendix C Experimental Details
@@
We use three well established datasets; a green fluorescent protein (GFP) from Aequorea Victoria (Sarkisyan et al., 2016), an adeno-associated virus (AAV) Bryant et al. (2021); and DNA binding activity to a human transcription factor (TFBIND8) (Trabucco et al., 2022; Barrera et al., 2016). These datasets have been used variously by Brookes & Listgarten (2018); Brookes et al. (2019); Angermueller et al. (2019); Kirjner et al. (2024); Jain et al. (2022) among others. The GFP task is to maximize fluorescence, this protein consists of 238 amino acids, of which 237 can mutate. The AAV task us to maximize the genetic payload that can be delivered, and the associated protein has 28 amino acids, all of which can mutate. A complete combinatorial assessment is infeasible for these tasks, and so we use the convolution neural network oracle presented in Kirjner et al. (2024) as in-silico ground truth. TFBIND8 contains a complete combinatorial assessment of the effect of changing 8 nucleotides on binding to human transcription factor SIX6 REF R1 (Barrera et al., 2016). The dataset we use contains all 65536 sequences prepared by Trabucco et al. (2022).
We also use two novel datasets from recent works that experimentally assess the (near) complete combinatorial space of short sequences. The first dataset measures the antibiotic resistance of Escherichia coli metabolic gene folA, which encodes dihydrofolate reductase (DHFR) (Papkou et al., 2023). Only a sub-sequence of this gene is varied (9 nucleic acids which encode 3 amino acids), and so a near-complete (99.7%) combinatorial scan is available. For variants that have no fitness (resistance) data available, we give a score of $-1$. The next dataset is near-complete combinatorial scan of four interacting amino acid residues near the active site of the enzyme tryptophan synthase (TrpB) (Johnston et al., 2024), with 159,129 unique sequences and fitness values, we use $-0.2$ for the missing fitness values (we do not use the authors’ imputed values). These residues are explicitly shown to exhibit epistasis – or non-additive effects on catalytic function – which makes navigating this landscape a more interesting challenge from an optimization perspective.
Finally, we use the recently proposed Ehrlich functions (Stanton et al., 2024) benchmark. These functions are challenging closed form biological analogues, specifically designed to test BBO methods on high dimensional sequence design tasks without having to resort to physical experimentation or machine learning oracles. We use the poli and poli-baselines software package for the benchmark and baselines (González-Duque et al., 2024), and test on both the original holo implementation (Stanton et al., 2024) as well as the native poli implementation of these functions.
The properties of these datasets and benchmarks are presented in Table C.1.
@@
Ehrlich-15 20 15 $20^{15}$ $20^{15}$
Ehrlich-32 20 32 $20^{32}$ $20^{32}$
Ehrlich-64 20 64 $20^{64}$ $20^{64}$
Table C.1: Alphabet size, sequence length, and number of available sequences for each of the datasets we use in this work.
We optimize VSD, CbAS, DbAS and BORE for a minimum of 3000 iterations each round (5000 for all experiments but the Ehrlich functions) using Adam (Kingma & Ba, 2014). When we use a CPE, AdaLead’s $\kappa$ parameter is set to 0.5 since the CPE already incorporates the appropriate threshold.
C.1 Fitness Landscapes Settings
For the DHFR and TrpB experiments we set maximum fitness in the training dataset to be that of the wild type, and $\tau$ to be slightly below the wild type fitness value (so we have $\sim\!10$ positive examples to train the CPE with). We use a randomly selected $N_{\textrm{train}}=2000$ below the wild-type fitness to initially train the CPE, we also explicitly include the wild-type. The thresholds and wild-type fitness values are; DHRF: $\tau=-0.1$, $y_{\textrm{wt}}=0$, TrpB: $\tau=0.35$, $y_{\textrm{wt}}=0.409$. We follow the same procedure for the TFBIND8 experiment, however, there is no notion of a wild-type sequence in this data, and so we set $\tau=0.75$, and $y_{\textrm{train max}}=0.85$. We use a uniform prior over sequences, $p({\mathbf{x}})=\prod^{M}_{m=1}\mathrm{Categ}({x_{m}}|{\mathbf{1}\cdot|\mathcal{V}|^{-1}})$, since these are relatively small search spaces, and the sub-sequences of nucleic/amino acids have been specifically selected for their task. Similarly, we find that relatively simple independent (mean-field) variational distributions of the form in Equation 20 and MLP based CPE s work best for these experiments (details in Sec. C.4).
C.2 Black-box Optimization Settings
We follow Kirjner et al. (2024) in the experimental settings for the AAV and GFP datasets, but we modify the maximum fitness training point and training dataset sizes to make them more amenable to a sequential optimization setting. The initial percentiles, schedule, and max training fitness values are; AAV: $p_{0}=0.8$, $\eta=0.7$, $y_{\textrm{max}}=5$, GFP: $p_{0}=0.8$, $\eta=0.7$ $y_{\textrm{max}}=1.9$. We aim for $p_{T}=0.99$. The edit distance between $\mathbf{x}^{*}$ and the fittest sequence in the CPE training data is 8 for GFP, and 13 for AAV. We again use a random $N_{\textrm{train}}=2000$ for training the CPE s, which in this case are CNNs – architecture specifics are in Sec. C.4.
For the Ehrlich function experiment, we use sequence lengths of $M=\{15,32,64\}$ with 2 motifs for the shorter sequence lengths, and 8 motifs for $M=64$. All use a motif length of 4 and a quantization of 4. $B=128$, $T=32$ and only 128 random samples of the function are used for $\mathcal{D}_{N}$ – these are resampled for each seed. As before, 5 different random seeds are used for these trials, and for VSD we use an the same scheduling function for $\tau_{t}$ as in Equation 18, with $p_{0}=0.5$ and $\eta=0.87$ (so $p_{T}=0.99$). The lower initial percentile is used since the training dataset is much smaller than in the other experiments, and we find allowing for more exploration initially improves VSD’s performance.
In these higher dimensional settings, we find that performance of the methods heavily relies on using an informed prior (in the case of VSD and CbAS), or initial variational distribution (in the case of DbAS and BORE). To this end, we follow Brookes et al. (2019) and fit the initial variational distribution to the CPE training sequences (regardless of fitness), but we use maximum likelihood. For the more complex variational distributions (LSTM and transformer), we have to be careful not to over-fit – so we implement early stopping and data augmentation techniques. Then for VSD and CbAS we copy this distribution and fix its parameters for the remainder of the experiment for use as a prior. We also use this prior for the Random method, but AdaLead and PEX use alternative generative heuristics. For these experiments we use the simple independent variational distribution and the same LSTM and causal decoder-only transformer models from Sec. 4.1.
C.3 Variational Distributions
In this section we summarize the main variational distribution architectures considered for VSD, BORE, CbAS and DbAS, and the sampling distributions for the Random baseline method. Somewhat surprisingly, we find that we often obtain good results for the biological sequence experiments using a simple independent (or mean-field) variational distribution, especially in lower dimensional settings,